1.How much charge does a particle have if it's electric field intensity at a point 10.0cm away is 5.20x10³N/C?

2.The primary winding of a transformer has 1000 turns while it's secondary winding has 500 turns.

A. If the input voltage is 220V, what is the output voltage?

B. Suppose that the current in the secondary is winding is 0.5A, what current flows in the primary winding?



1
Expert's answer
2022-05-15T17:02:29-0400

1.

"E=k\frac{q}{r^2}"

"q=\frac{Er^2}{k}"

"q=\frac{5.20*10^3*0.10^2}{9*10^9}=5.78*10^{-9}\rm C"

2.

"V_2=V_1\frac{N_2}{N_1}"

"V_2=220*\frac{500}{1000}=110\:\rm V"